OpenStudy (anonymous):

*fan and medal awarded* *warning: it looks long but its not, only 2 parts* Part 1: For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 9 3x + 4y = 26

OpenStudy (anonymous):

i found this y=11-2x 3x+4(11-2x)=24 3x+44-8x=24 -5x=-20 x=4 y=11-8 y=3 lollipop costs 4dollars, cola 3 dollars

OpenStudy (anonymous):

part 2 is right

OpenStudy (anonymous):

okay, now i jus need part 1

OpenStudy (anonymous):

wait im confused. whats part 1 and what is part 2

OpenStudy (anonymous):

Sum of twice a number and another is 9, if you triple the first and multiply the second by 4 then they add to 26 find the numbers. 2x + y = 9 3x + 4y = 26

OpenStudy (anonymous):

wait im confused. whats part 1 and what is part 2

OpenStudy (anonymous):

this is part 2

OpenStudy (anonymous):

okay thanks. can u fully explain it to me ?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

@begzat